If a bank's total reserves are $13.1bln., reserve requirement ratio is 5.9%, and the bank's total deposits are $112.5bln., what are the bank's excess reserves, in $bln (to the nearest $0.01bln)?
Note: Excess reserves may be negative.
Answer #1

Bank Deposits = $112.5 b

Reserve Requirement = 5.9% of Bank Deposits = 0.059*112.5 = $6.64 b

Total Reserves = $13.1 b

Hence, Excess Reserves = Total Reserve - Reserve Requirement = 13.1 - 6.64 = $6.46 b